# Fermi Machine — Quantum Many-Body Solver Derived from Correspondence between Noninteracting and Strongly Correlated Fermions

### v1

@@ Abstract
Stimulated by the successful descriptions of strongly correlated electron systems by fractionalized fermions, correspondence between interacting fermions and non-interacting multi-component fermions is formulated in examples of the Hubbard model. The formalism enables constructions of the neural network for a quantum many-body solver represented by coupled noninteracting fermions. After showing the exact correspondence of 1- and 2-site Hubbard model to two-component noninteracting fermions, numerical algorithm of the quantum machine learning for the Hubbard model is proposed. Benchmark for the 4-site systems is successfully presented and promising future directions as well as implications are discussed.
@@ 1 Introduction
Meanwhile, numerical method to construct ground state wave functions of strongly interacting quantum many-body systems is one of the grand challenges in physics. Various algorithms have been proposed such as a variety of auxiliary-field quantum Monte Carlo [14, 15, 16], variational Monte Carlo [17, 18, 19, 20, 21, 22], density matrix renormalization group [23], tensor network [24, 25, 26], density matrix embedding [27] and dynamical mean-field theory [28, 29]. Recently, neural network with Boltzmann machine has established an accurate way to approximate ground states of quantum spin systems [30] as well as itinerant fermion systems [31]. Vision transformer, another architecture of the neural network, has also shown the state-of-the-art accuracy for quantum spin models [32, 33]. We note that these neural networks are constructed by introducing hidden variables that are described by classical degrees of freedom.
In this paper, we discuss mapping of the Hubbard model (or in more general interacting lattice fermions) to noninteracting multi-component fermion models in the ground states that materializes the fractionalization, from which we propose a quantum machine learning algorithm aiming at an efficient quantum many-body solver. Here, we introduce a fermion system as the hidden part coupled to the physical system, instead of the classical one such as the Ising spins in the previous neural networks.
In Secs. I and II, the equivalence between the Hubbard model and the two-component fermion model (TCFM) is shown for the atomic limit and for the 2-site systems, respectively. Based on this equivalence, we propose a quantum machine learning algorithm in Sec. III, where the Hubbard model is approximated by a deep-layer representation of noninteracting multi-component fermion model. The couplings between the layers are represented by the hybridization of fermions belonging to neighboring layers. This method is viewed as an extension of restricted Boltzmann-machine representation of the neural quantum states proposed by Carleo and Troyer [30] to a quantum neural network, where the Ising hidden variable in the Boltzmann machine is replaced by the fermionic operators that allows hybridization to generate quantum entanglement. Furthermore, it is also regarded as an embodiment of the electron fractionalization [9, 10], which has succeeded in solving a number of experimental puzzles in the cuprate high-$T_{c}$ superconductors [8, 12, 10, 13]. This quantum neural network can also be viewed as a tractable realization of the ground states by a systematic expansion to take account of correlation effects similarly to the formulation by the equation of motion method or equivalently the continued fraction expansion of the Green’s function for the excited states [34, 35, 36]. The present method does not seem to have difficulty of exponentially increasing terms encountered in the conventional equation of motion method.
@@ 2 Hubbard model in the atomic limit
In this section, the result shown in Appendix A of Ref. [9] is briefly summarized for the self-contained presentation. We consider the Hubbard $U$ term
@@
with $n_{\sigma}=c^{\dagger}_{\sigma}c_{\sigma}$ for the creation (annihilation) operatores $c_{\sigma}^{\dagger}$ ($c_{\sigma}$) of the spin $\sigma$, and introduce an auxiliary fermion
@@
The bonding and antibonding states given by Eqs. (17) and (18), respectively, are nothing but the operators for the lower and upper Hubbard levels, respectively, in the Hubbard model whose averaged energies are given by $E=0$ and $U$. Namely, aside from the normalization factor, ${b}_{\sigma}^{\dagger}$ creates an electron with the spin $\sigma$ when it is not occupied by the opposite spin electron (namely it fills the lower Hubbard state in the corresponding Hubbard model) as is apparent from the last equation of Eq.(17). On the other hand, $a_{\sigma}^{\dagger}$, creates an electron with the spin $\sigma$, when the opposite spin electron already exists, namely it creates a doublon at the upper Hubbard state in the Hubbard model. It helps an intuitive interpretation of the correspondence between the Hubbard model and the TCFM. With the choice Eq.(8), $\mu_{a}=0$ and $\mu_{b}=U$ are obtained and the Hubbard gap is reproduced. In addition, the ground state $b_{\sigma}^{\dagger}|0\rangle$ indeed satisfies the form of Eq.(13). Then the mapping between Eqs.(1) and (4) becomes exact. In this correspondence, the Mott gap in the Hubbard model is interpreted as the hybridization gap in the TCFM.
@@
$G_{\tilde{c}_{\sigma},\tilde{d}_{\sigma}^{\dagger}}(\omega)$ $=$ $G_{\tilde{d}_{\sigma},\tilde{c}_{\sigma}^{\dagger}}(\omega)=\frac{-\Lambda}{(\omega-\mu_{\tilde{c}})(\omega-\mu_{\tilde{d}})-\Lambda^{2}}$
$G_{\tilde{d}_{\sigma},\tilde{d}_{\sigma}^{\dagger}}(\omega)$ $=$ $\frac{1}{\omega-\mu_{\tilde{d}}-\frac{\Lambda^{2}}{\omega-\mu_{\tilde{c}}}},$ (19)
@@
$=$ $\frac{1}{2}\left[\frac{1}{\omega}+\frac{1}{\omega-U}\right]$ (21)
@@ 3 Two-site Hubbard model
${\cal H}$ $=$ ${\cal H}_{t}+{\cal H}_{U}$ (23)
${\cal H}_{t}$ $=$ $\sum_{\sigma}[-t(c_{1\sigma}^{\dagger}c_{2\sigma}+c_{2\sigma}^{\dagger}c_{1\sigma})+\mu\sum_{i=1,2}n_{i,\sigma}]$ (24)
${\cal H}_{U}$ $=$ $U\sum_{i=1,2}n_{i\uparrow}n_{i\downarrow}$ (25)
@@
${\cal H}=\left(\begin{array}[]{c|cccc}&|\uparrow,\downarrow\rangle&|\downarrow,\uparrow\rangle&|\uparrow\downarrow,0\rangle&|0,\uparrow\downarrow\rangle\\ \hline\cr\langle\uparrow,\downarrow|&0&0&-t&-t\\ \langle\downarrow,\uparrow|&0&0&-t&-t\\ \langle\uparrow\downarrow,0|&-t&-t&U&0\\ \langle 0,\uparrow\downarrow|&-t&-t&0&U\end{array}\right)$ (31)
@@
$\frac{1}{\sqrt{2(R^{2}+Q^{2})}}(R|\uparrow,\downarrow\rangle+R|\downarrow,\uparrow\rangle-Q|\uparrow\downarrow,0\rangle-Q|0,\uparrow\downarrow\rangle)$ (34)
@@
$-(R(\tilde{b}_{1,\uparrow}^{\dagger}+\tilde{b}_{2,\uparrow}^{\dagger})+Q(\tilde{a}_{1,\uparrow}^{\dagger}+\tilde{a}_{2,\uparrow}^{\dagger}))$
@@
$\Leftrightarrow$ ${\rm Eq.(29)}$
Here, in the Hubbard model, we used the relations $(\tilde{c}_{i,\sigma}^{\dagger}-\tilde{d}_{i,\sigma}^{\dagger})|0\rangle=0$, $(\tilde{c}_{i,\sigma}^{\dagger}+\tilde{d}_{i,\sigma}^{\dagger})(\tilde{c}_{i,-\sigma}^{\dagger}+\tilde{d}_{i,-\sigma}^{\dagger})|0\rangle=0$ and $(\tilde{c}_{1i,\sigma}^{\dagger}-\tilde{d}_{i,\sigma}^{\dagger})(\tilde{c}_{j,-\sigma}^{\dagger}+\tilde{d}_{j,-\sigma}^{\dagger})|0\rangle=0$ for $i\neq j$, because $\tilde{c}_{i,\sigma}^{\dagger}-\tilde{d}_{i,\sigma}^{\dagger}$ creates a doublon at the $i$th site singly occupied by a spin $-\sigma$ electron and $\tilde{c}_{i,\sigma}^{\dagger}+\tilde{d}_{i,\sigma}^{\dagger}$ creates an electron at the empty $i$th site. Mapping for the excited states are similarly shown. Therefore, the exact mapping of not only the ground state but also the whole structure of the spectra between the TCFM and Hubbard model is proven for the two-site system as well.
@@ 4 Fermi Machine as a Quantum Neural Network
We have shown that the Hubbard model has an exact mapping to the TCFM for one- and two-site problems and the ground state wave function of the Hubbard model can be constructed from the ground state of the TCFM. It demonstrates that the TCFM is able to describe the energy scales of the Mott gap separating the lower and upper Hubbard bands scaled by $E_{2}-E_{0}\propto U$ and the singlet-triplet gap ($E_{1}-E_{0}$ and $E_{3}-E_{2}$) associated with the superexchange interaction $J$, which scales to $4t^{2}/U$ in the strong coupling limit. Up to the 2 sites, the exact ground and excited states of the Hubbard model are represented by the corresponding ground and excited states of the TCFM, respectively by optimizing the variational parameters given by Eqs.(8) and (39), namely, $\mu_{\tilde{c}},\mu_{\tilde{d}},\Lambda,t_{\tilde{c}}$ and $t_{\tilde{d}}$.
@@
${\cal H}$ $=$ ${\cal H}_{t}+{\cal H}_{U}$ (50)
${\cal H}_{t}$ $=$ $\sum_{\langle i,j\rangle,\sigma}[-t(c_{i\sigma}^{\dagger}c_{j\sigma}+{\rm H.c.}+\mu\sum_{i}^{N_{s}}n_{i,\sigma}]$ (51)
@@
$G(q,\omega)$ $=$ $\frac{1}{\omega-\epsilon_{c}(q)-\Sigma_{1}(q,\omega)}$ (53)
$\Sigma_{1}(q,\omega)$ $=$ $\frac{\eta_{1}(q)}{\omega-\epsilon_{1}(q)-\Sigma_{2}(q,\omega)}$
@@
$\eta_{1}(q)$ $=$ $\Lambda$
$\epsilon_{m}(q)$ $=$ $\frac{1}{N_{s}}\sum_{\langle\ell,j\rangle}e^{iq\ell-j)}(-t_{\tilde{d},\ell,j}^{(m)})+\mu_{\tilde{d},i}^{(m)}$
@@
Figure 1: Illustration of the architecture of Fermi machine network based on the MCFM compared to the Hubbard model.
@@
$|\Psi_{0}^{\rm MCFM}\rangle$ $=$ $\begin{pmatrix}|\Psi_{0\ \uparrow}^{\rm MCFM}\rangle&0\\ 0&|\Psi_{0\ \downarrow}^{\rm MCFM}\rangle\end{pmatrix}$ (64)
@@
$|\Psi_{0}^{\rm Hub}\rangle$ $=$ $\begin{pmatrix}|\Psi_{0\ \uparrow}^{\rm Hub}\rangle&0\\ 0&|\Psi_{0\ \downarrow}^{\rm Hub}\rangle\end{pmatrix}$ (67)
@@
$f(\beta,q)$ $=$ $\frac{1}{1+e^{\beta{\cal E}(q)}}$ (76)
@@
where $\beta$ is an additional variational parameter, $|\Psi_{\sigma}^{\rm MCFM}(q)\rangle$ is the eigenstate of Eq.(55) with the momentum $q$ and $|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ is the same as Eq.(70). The product with respect to $q$ is over the full Brillouin zone for Eq. (75). If we take $\beta\rightarrow\infty$, it is reduced to the original algorithm using Eqs.(69) and (70). Therefore this Fermi machine extension certainly enhances the representability with a strategy conceptually similar to the Boltzmann machine [30, 31] but by using substantially different quantum algorithm.
@@
if the $i$th spin-${\sigma}$ electron is at the site $n$ in the sample.
@@
In the matrix representation, $|\Psi_{\sigma}^{\rm MCFM}(q)\rangle$ and $|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ are $N_{s}$-component vectors and for instance, $\langle\Psi_{\sigma}^{\rm MCFM}(q)|{\cal H}^{\rm MCFM}|\Psi_{\sigma}^{\rm MCFM}(q)\rangle$ is a scalar for diagonalized $N_{s}\times N_{s}$ matrix of ${\cal H}^{\rm MCFM}$.
@@
${\cal H}_{\rm nonlocal}$ $=$ ${\cal H}^{(\tilde{d})}_{\rm nonlocal}+{\cal H}^{(\tilde{c}\tilde{d})}_{\rm nonlocal}$ (83)
@@
$+$ $\sum_{i,\sigma}\mu_{\tilde{d},i}^{(M+1)}n_{\tilde{d},i,\sigma}^{(M+1)}$
@@
$+$ $\alpha_{M+1,q}\sum_{\ell,\ell^{\prime},\sigma^{\prime}}\frac{e^{iq\ell^{\prime}}}{\sqrt{N_{s}}}\Theta_{\sigma^{\prime},\sigma}(\ell^{\prime}-\ell)c_{\ell,\sigma}^{\dagger}(-)^{n_{\ell^{\prime},\sigma^{\prime}}}]|0\rangle$
which means that the spin dependent $\Theta_{\sigma,\sigma^{\prime}}(\ell^{\prime}-\ell)$ contains the spin parallel ($\sigma=\sigma^{\prime}$) and antiparallel ($\sigma=-\sigma^{\prime}$) components. $\Theta_{\sigma,\sigma^{\prime}}(\ell^{\prime}-\ell)$ is regarded as additional variational parameters. The present extension explicitly allows dependence of the weight of the wave function at the $\ell$th site with spin $\sigma$ on the occupation of parallel ($\sigma$) or antiparallel ($-\sigma$) component of the fermion at the $\ell^{\prime}$th site, which plays a role similar to Jastrow factor in the conventional variational Monte Carlo or nonlocal coupling between the visible (physical) and hidden variables in the Boltzmann machine. However, the present nonlocal hybridization explicitly introduces spatially extended quantum entanglement between up- and down-spin electrons. In addition, the present scheme does not break the SU(2) symmetry in contrast to the spin Jastrow factor in the variational Monte Carlo and the Boltzmann machine. This is because the amplitude of $|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ for the component of the up spin at the $i$th site and the down spin at the $j$th site has the same value with the opposite configuration and the matrix element of spin diagonal component $\langle\Psi_{\sigma}^{\rm Hub}(q)|S_{i}^{z}S_{j}^{z}|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ becomes the same as the off-diagonal one $\langle\Psi_{\sigma}^{\rm Hub}(q)|S_{i}^{+}S_{j}^{-}|\Psi_{\sigma}^{\rm Hub}(q)\rangle=\langle\Psi_{\sigma}^{\rm Hub}(q)|S_{i}^{-}S_{j}^{+}|\Psi_{\sigma}^{\rm Hub}(q)\rangle$. This nonlocal hybridization part may also be extended to deeper layers of $M+3$rd, $M+4$th $\cdots$ up to $M+M^{\prime}$th layer if necessary. In this extension, Eqs. (62) and (63) become more complicated, because of the nonlocality but may enable the orthogonality of $\tilde{d}_{i,\sigma}^{(m)}$ in the Hubbard terminology expected for Eq.(53). Then, if we include the $k$-independent component of $\Lambda$ in Eq.(84), the local part represented by the first term of the r.h.s. of Eq.(LABEL:Psi_Hub_qsigma2) can be absorbed, but still only the antiparallel spin combination is relevant for that $k$-independent component. In this sense, further generalized simpler architecture would be represented by the following unified form instead of Eqs. (70) and (LABEL:Psi_Hub_qsigma2):
$|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ $=$ $\sum_{m=0}^{M}\alpha_{m,q}\sum_{\ell,\ell^{\prime},\sigma^{\prime}}\frac{e^{iq\ell^{\prime}}}{\sqrt{N_{s}}}\Theta_{\sigma^{\prime},\sigma}^{m}(\ell^{\prime}-\ell)c_{\ell,\sigma}^{\dagger}$ (86)
$\times$ $(-)^{mn_{\ell^{\prime},\sigma^{\prime}}}|0\rangle.$

### v2

@@ Abstract
Stimulated by the successful descriptions of strongly correlated electron systems by fractionalized fermions, correspondence between interacting fermions and non-interacting multi-component fermions is formulated in examples of the Hubbard model. The formalism enables constructions of the neural network for quantum many-body solvers represented by coupled noninteracting fermions. After showing the exact correspondences of 1- and 2-site Hubbard models to two-component noninteracting fermions, a numerical algorithm of the quantum machine learning for the Hubbard model is proposed. Benchmark for the 4-site systems is successfully presented and promising future directions as well as implications are discussed.
@@ 1 Introduction
Meanwhile, numerical method to construct ground state wave functions of strongly interacting quantum many-body systems is one of the grand challenges in physics. Various algorithms such as a variety of auxiliary-field quantum Monte Carlo [14, 15, 16], variational Monte Carlo [17, 18, 19, 20, 21, 22], density matrix renormalization group [23], tensor network [24, 25, 26], density matrix embedding [27] and dynamical mean-field theory [28, 29] have been proposed. Recently, neural network with Boltzmann machine has established an accurate way to approximate ground states of quantum spin systems [30] as well as itinerant fermion systems [31]. Vision transformer, another architecture of the neural network, has also shown the state-of-the-art accuracy for quantum spin models [32, 33]. We note that these neural networks are constructed by introducing hidden variables that are described by classical degrees of freedom.
In this paper, we discuss mapping of the Hubbard model (or in more general, interacting lattice fermions) to noninteracting multi-component fermion models in the ground states that materializes the fractionalization, from which we propose a quantum machine learning algorithm aiming at an efficient quantum many-body solver. Here, we introduce a fermion system as the hidden part coupled to the physical system, instead of the classical one such as the Ising spins in the previous neural networks.
In Secs. I and II, the equivalence between the Hubbard model and the two-component fermion model (TCFM) is shown for the atomic limit and for the 2-site systems, respectively. Based on this equivalence, we propose a quantum machine learning algorithm in Sec. III, where the Hubbard model is approximated by a deep-layer representation of noninteracting multi-component fermion models. The couplings between the layers are represented by the hybridization of fermions belonging to neighboring layers. This method is viewed as an extension of restricted Boltzmann-machine representation of the neural quantum states proposed by Carleo and Troyer [30] to a quantum neural network, where the Ising hidden variable in the Boltzmann machine is replaced by the fermionic operators that allows hybridization to generate quantum entanglement. Furthermore, it is also regarded as an embodiment of the electron fractionalization [9, 10], which has succeeded in solving a number of experimental puzzles in the cuprate high-$T_{c}$ superconductors [8, 12, 10, 13]. This quantum neural network can also be viewed as a tractable realization of the ground states by a systematic expansion to take account of correlation effects similarly to the formulation by the equation of motion method or equivalently the continued fraction expansion of the Green’s function for the excited states [34, 35, 36]. The present method does not seem to have difficulty of exponentially increasing terms encountered in the conventional equation of motion method.
@@ 2 Hubbard model in the atomic limit
[…]
@@
with $n_{\sigma}=c^{\dagger}_{\sigma}c_{\sigma}$ for the creation (annihilation) operators $c_{\sigma}^{\dagger}$ ($c_{\sigma}$) of the spin $\sigma$, and introduce an auxiliary fermion
@@
The bonding and antibonding states are given by Eqs. (17) and (18), respectively, which are nothing but the operators for the lower and upper Hubbard levels, respectively, in the Hubbard model whose averaged energies are given by $E=0$ and $U$. Namely, aside from the normalization factor, ${b}_{\sigma}^{\dagger}$ creates an electron with the spin $\sigma$ when it is not occupied by the opposite spin electron (namely it fills the lower Hubbard state in the corresponding Hubbard model) as is apparent from the last equation of Eq.(17). On the other hand, $a_{\sigma}^{\dagger}$, creates an electron with the spin $\sigma$, when the opposite spin electron already exists, namely it creates a doublon at the upper Hubbard state in the Hubbard model. It helps an intuitive interpretation of the correspondence between the Hubbard model and the TCFM. With the choice Eq.(8), $\mu_{a}=0$ and $\mu_{b}=U$ are obtained and the Hubbard gap is reproduced. In addition, the ground state $b_{\sigma}^{\dagger}|0\rangle$ indeed satisfies the form of Eq.(13). Then the mapping between Eqs.(1) and (4) becomes exact. In this correspondence, the Mott gap in the Hubbard model is interpreted as the hybridization gap in the TCFM.
@@
$G_{\tilde{c}_{\sigma},\tilde{d}_{\sigma}^{\dagger}}(\omega)$ $=$ $G_{\tilde{d}_{\sigma},\tilde{c}_{\sigma}^{\dagger}}(\omega)=\frac{-\Lambda}{(\omega-\mu_{\tilde{c}})(\omega-\mu_{\tilde{d}})-\Lambda^{2}},$
$G_{\tilde{d}_{\sigma},\tilde{d}_{\sigma}^{\dagger}}(\omega)$ $=$ $\frac{1}{\omega-\mu_{\tilde{d}}-\frac{\Lambda^{2}}{\omega-\mu_{\tilde{c}}}}.$ (19)
@@
$=$ $\frac{1}{2}\left[\frac{1}{\omega}+\frac{1}{\omega-U}\right].$ (21)
@@ 3 Two-site Hubbard model
${\cal H}$ $=$ ${\cal H}_{t}+{\cal H}_{U},$ (23)
${\cal H}_{t}$ $=$ $\sum_{\sigma}[-t(c_{1\sigma}^{\dagger}c_{2\sigma}+c_{2\sigma}^{\dagger}c_{1\sigma})+\mu\sum_{i=1,2}n_{i,\sigma}],$ (24)
${\cal H}_{U}$ $=$ $U\sum_{i=1,2}n_{i\uparrow}n_{i\downarrow}.$ (25)
@@
${\cal H}=\left(\begin{array}[]{c|cccc}&|\uparrow,\downarrow\rangle&|\downarrow,\uparrow\rangle&|\uparrow\downarrow,0\rangle&|0,\uparrow\downarrow\rangle\\ \hline\cr\langle\uparrow,\downarrow|&0&0&-t&-t\\ \langle\downarrow,\uparrow|&0&0&-t&-t\\ \langle\uparrow\downarrow,0|&-t&-t&U&0\\ \langle 0,\uparrow\downarrow|&-t&-t&0&U\end{array}\right).$ (31)
@@
$\frac{1}{\sqrt{2(R^{2}+Q^{2})}}(R|\uparrow,\downarrow\rangle+R|\downarrow,\uparrow\rangle-Q|\uparrow\downarrow,0\rangle-Q|0,\uparrow\downarrow\rangle).$ (34)
@@
-- $(R(\tilde{b}_{1,\uparrow}^{\dagger}+\tilde{b}_{2,\uparrow}^{\dagger})+Q(\tilde{a}_{1,\uparrow}^{\dagger}+\tilde{a}_{2,\uparrow}^{\dagger}))$
@@
$\Leftrightarrow$ ${\rm Eq.(29)}.$
Here, in the Hubbard model, we used the relations $(\tilde{c}_{i,\sigma}^{\dagger}-\tilde{d}_{i,\sigma}^{\dagger})|0\rangle=0$, $(\tilde{c}_{i,\sigma}^{\dagger}+\tilde{d}_{i,\sigma}^{\dagger})(\tilde{c}_{i,-\sigma}^{\dagger}+\tilde{d}_{i,-\sigma}^{\dagger})|0\rangle=0$ and $(\tilde{c}_{1i,\sigma}^{\dagger}-\tilde{d}_{i,\sigma}^{\dagger})(\tilde{c}_{j,-\sigma}^{\dagger}+\tilde{d}_{j,-\sigma}^{\dagger})|0\rangle=0$ for $i\neq j$, because $\tilde{c}_{i,\sigma}^{\dagger}-\tilde{d}_{i,\sigma}^{\dagger}$ creates a doublon at the $i$th site singly occupied by a spin $-\sigma$ electron and $\tilde{c}_{i,\sigma}^{\dagger}+\tilde{d}_{i,\sigma}^{\dagger}$ creates an electron at the empty $i$th site. Mapping for the excited states is similarly shown. Therefore, the exact mapping of not only the ground state but also the whole structure of the spectra between the TCFM and Hubbard model is proven for the two-site system as well.
@@ 4 Fermi Machine as a Quantum Neural Network
We have shown that the Hubbard model has an exact mapping to the TCFM for one- and two-site problems and the ground-state wave function of the Hubbard model can be constructed from the ground state of the TCFM. It demonstrates that the TCFM is able to describe the energy scales of the Mott gap separating the lower and upper Hubbard bands scaled by $E_{2}-E_{0}\propto U$ and the singlet-triplet gap ($E_{1}-E_{0}$ and $E_{3}-E_{2}$) associated with the superexchange interaction $J$, which scales to $4t^{2}/U$ in the strong coupling limit. Up to the 2 sites, the exact ground and excited states of the Hubbard model are represented by the corresponding ground and excited states of the TCFM, respectively by optimizing the variational parameters given by Eqs.(8) and (39), namely, $\mu_{\tilde{c}},\mu_{\tilde{d}},\Lambda,t_{\tilde{c}}$ and $t_{\tilde{d}}$.
@@
${\cal H}$ $=$ ${\cal H}_{t}+{\cal H}_{U},$ (50)
${\cal H}_{t}$ $=$ $\sum_{\langle i,j\rangle,\sigma}[-t(c_{i\sigma}^{\dagger}c_{j\sigma}+{\rm H.c.}+\mu\sum_{i}^{N_{s}}n_{i,\sigma}],$ (51)
@@
$G(q,\omega)$ $=$ $\frac{1}{\omega-\epsilon_{c}(q)-\Sigma_{1}(q,\omega)},$ (53)
$\Sigma_{1}(q,\omega)$ $=$ $\frac{\eta_{1}(q)}{\omega-\epsilon_{1}(q)-\Sigma_{2}(q,\omega)},$
@@
$\eta_{1}(q)$ $=$ $\Lambda,$
$\epsilon_{m}(q)$ $=$ $\frac{1}{N_{s}}\sum_{\langle\ell,j\rangle}e^{iq\ell-j)}(-t_{\tilde{d},\ell,j}^{(m)})+\mu_{\tilde{d},i}^{(m)},$
@@
Figure 1: (Color online): Illustration of the architecture of Fermi machine network based on the MCFM compared to the Hubbard model.
@@
$|\Psi_{0}^{\rm MCFM}\rangle$ $=$ $\begin{pmatrix}|\Psi_{0\ \uparrow}^{\rm MCFM}\rangle&0\\ 0&|\Psi_{0\ \downarrow}^{\rm MCFM}\rangle\end{pmatrix},$ (64)
@@
$|\Psi_{0}^{\rm Hub}\rangle$ $=$ $\begin{pmatrix}|\Psi_{0\ \uparrow}^{\rm Hub}\rangle&0\\ 0&|\Psi_{0\ \downarrow}^{\rm Hub}\rangle\end{pmatrix},$ (67)
@@
$f(\beta,q)$ $=$ $\frac{1}{1+e^{\beta{\cal E}(q)}},$ (76)
@@
where $\beta$ is an additional variational parameter, $|\Psi_{\sigma}^{\rm MCFM}(q)\rangle$ is the eigenstate of Eq.(55) with the momentum $q$ and $|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ is the same as Eq.(70). The product with respect to $q$ is over the full Brillouin zone for Eq. (75). If we take $\beta\rightarrow\infty$, it is reduced to the original algorithm using Eqs.(69) and (70). Therefore this Fermi machine extension certainly enhances the representability with a strategy conceptually similar to the Boltzmann machine [30, 31] but by using a substantially different quantum algorithm.
@@
if the $i$th spin-${\sigma}$ electron is at the site $n$ in the given sample.
@@
In the matrix representation, $|\Psi_{\sigma}^{\rm MCFM}(q)\rangle$ and $|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ are $N_{s}$-component vectors and for instance, $\langle\Psi_{\sigma}^{\rm MCFM}(q)|{\cal H}^{\rm MCFM}|\Psi_{\sigma}^{\rm MCFM}(q)\rangle$ is a scalar for diagonalized $N_{s}\times N_{s}$ matrix of ${\cal H}^{\rm MCFM}$ in the momentum representation of Eqs. (55)-(58).
@@
${\cal H}_{\rm nonlocal}$ $=$ ${\cal H}^{(\tilde{d})}_{\rm nonlocal}+{\cal H}^{(\tilde{c}\tilde{d})}_{\rm nonlocal},$ (83)
@@
$+$ $\sum_{i,\sigma}\mu_{\tilde{d},i}^{(M+1)}n_{\tilde{d},i,\sigma}^{(M+1)},$
@@
$+$ $\alpha_{M+1,q}\sum_{\ell,\ell^{\prime},\sigma^{\prime}}\frac{e^{iq\ell^{\prime}}}{\sqrt{N_{s}}}\Theta_{\sigma^{\prime},\sigma}(\ell^{\prime}-\ell)c_{\ell,\sigma}^{\dagger}(-)^{n_{\ell^{\prime},\sigma^{\prime}}}]|0\rangle,$
which means that the spin dependent $\Theta_{\sigma,\sigma^{\prime}}(\ell^{\prime}-\ell)$ contains the spin parallel ($\sigma=\sigma^{\prime}$) and antiparallel ($\sigma=-\sigma^{\prime}$) components. $\Theta_{\sigma,\sigma^{\prime}}(\ell^{\prime}-\ell)$ is regarded as additional variational parameters. The present extension explicitly allows dependence of the weight of the wave function at the $\ell$th site with spin $\sigma$ on the occupation of parallel ($\sigma$) or antiparallel ($-\sigma$) component of the fermion at the $\ell^{\prime}$th site, which plays a role similar to Jastrow factor in the conventional variational Monte Carlo or nonlocal coupling between the visible (physical) and hidden variables in the Boltzmann machine. However, the present nonlocal hybridization explicitly introduces spatially extended quantum entanglement between up- and down-spin electrons. In addition, the present scheme does not break the SU(2) symmetry in contrast to the spin Jastrow factor in the variational Monte Carlo and the Boltzmann machine. This is because the amplitude of $|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ for the component of the up spin at the $i$th site and the down spin at the $j$th site has the same value with the opposite configuration and the matrix element of spin diagonal component $\langle\Psi_{\sigma}^{\rm Hub}(q)|S_{i}^{z}S_{j}^{z}|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ becomes the same as the off-diagonal one $\langle\Psi_{\sigma}^{\rm Hub}(q)|S_{i}^{+}S_{j}^{-}|\Psi_{\sigma}^{\rm Hub}(q)\rangle=\langle\Psi_{\sigma}^{\rm Hub}(q)|S_{i}^{-}S_{j}^{+}|\Psi_{\sigma}^{\rm Hub}(q)\rangle$. This nonlocal hybridization part may also be extended to deeper layers of $M+3$rd, $M+4$th $\cdots$ up to $M+M^{\prime}$th layer if necessary.
In this extension, Eqs. (62) and (63) become more complicated, because of the nonlocality but may enable the mutual orthogonality of $\tilde{d}_{i,\sigma}^{(m)}$ in the Hubbard terminology expected for Eq.(53). Then, if we include the $k$-independent component of $\Lambda$ in Eq. (84), the local part represented by the first term of the r.h.s. of Eq. (85) can be absorbed, but still only the antiparallel spin combination is relevant for that $k$-independent component. In this sense, further generalized simpler architecture would be represented by the following unified form instead of Eqs. (70) and (85):
$|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ $=$ $\sum_{m=0}^{M}\alpha_{m,q}\sum_{\ell,\ell^{\prime},\sigma^{\prime}}\frac{e^{iq\ell^{\prime}}}{\sqrt{N_{s}}}\Theta_{\sigma^{\prime},\sigma}^{(m)}(\ell^{\prime}-\ell)c_{\ell,\sigma}^{\dagger}$ (86)
×$\times(-)^{mn_{\ell^{\prime},\sigma^{\prime}}}|0\rangle.$